OpenStudy (anonymous):

here is a list of numbers: 0,3,5,7,12,29. find three numbers from the list with the range double the median.

hartnn (hartnn):

the median as of now is (5+7)/2 = 6 so range = 12. three numbers from the list with the range 12 ? sorry, i didn't get the question....

OpenStudy (shubhamsrg):

I guess 12 is the maximum value of the range ?

OpenStudy (anonymous):

i dont get it either i know the median and range

hartnn (hartnn):

range = maximum value in list - minimum value in the list.

OpenStudy (anonymous):

29-0

OpenStudy (anonymous):

?

OpenStudy (anonymous):

29-3?

hartnn (hartnn):

i think its range = maximum value in list - minimum value in the list+1. = 29-0 +1 = 30.

OpenStudy (anonymous):

not +1

OpenStudy (anonymous):

26

Parth (parthkohli):

No, just max - min.

hartnn (hartnn):

min value = 0 so, range = 29

OpenStudy (anonymous):

the range is definatly 24 and the median is 12

OpenStudy (anonymous):

0 is not a value

OpenStudy (anonymous):

\[0 \] is not a value

hartnn (hartnn):

median is the middle value, when the elements are in ascending/descending order. if there are even number of elements, then there will be 2 middle values and median = average of those 2 values, so median = (5+7)/2 =6 and 0 is considered as element and its min.

Parth (parthkohli):

But we need three numbers from this list, not the list itself.

hartnn (hartnn):

thats why i didn't get the question.....

Parth (parthkohli):

Obviously, \(0\) can't be in the list.

Parth (parthkohli):

But you can try 5,12 and 29...

OpenStudy (anonymous):

no it gives the infomation that range=24 and median=12 i understand how to calculate range and median and dont understand why it says that the abover is the range and median

OpenStudy (anonymous):

0 is in the list

Parth (parthkohli):

I mean, \(0\) can't be amongst the three numbers we wanna choose.

OpenStudy (anonymous):

oh yeah

hartnn (hartnn):

that..yes.

Parth (parthkohli):

We just have to choose three numbers from the list such that the three numbers have their median = 12 and range = 24.

Parth (parthkohli):

So observe,\[29 -5 = 24 \]

OpenStudy (anonymous):

how can there be a median and range of a single number

hartnn (hartnn):

then 5,12,29 is eligible list.

Parth (parthkohli):

So the middle element can be 12. And our set is 5,12,29

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

oh i get it

Parth (parthkohli):

5,12,29. The median is the middle number. Range = max - min.

hartnn (hartnn):

there can be a median and range of a single number, but its not the case here.

Parth (parthkohli):

we have to choose three numbers!

OpenStudy (anonymous):

what about the 3

Parth (parthkohli):

No, we can choose ANY three numbers.

OpenStudy (anonymous):

5 12 29?

Parth (parthkohli):

Yes

OpenStudy (anonymous):

but in the end there can only be 3 correct numbers

OpenStudy (anonymous):

what about 3

Parth (parthkohli):

Yes, and these are the three correct numbers... :|

OpenStudy (anonymous):

0 3 5 (emphasis on 3)

Parth (parthkohli):

Is 5 - 0 the double of 3?

OpenStudy (anonymous):

@ParthKohli i know

OpenStudy (anonymous):

no its 5

OpenStudy (anonymous):

3x2=6

OpenStudy (anonymous):

5-0 =5

Parth (parthkohli):

So 0,3,5 is not a valid list...

Parth (parthkohli):

The only valid list is 5,12,29. Deal with it!

OpenStudy (anonymous):

no whay did you not do 29-3

Parth (parthkohli):

Hmm?

hartnn (hartnn):

lol :P because 0 is minimum .

OpenStudy (anonymous):

0 is not a value is it but its got nothing to do with 3

hartnn (hartnn):

0 is a value, valid element of list and minimum element of this list so range = 29-0 = 29

Parth (parthkohli):

No, my point was that 0 could not have been in the three numbers because in that case, we'd wanna have something in the form \(0,n,2n\). There are no two numbers in the form \(n\) and \(2n\).

Parth (parthkohli):

In the list, that is.

OpenStudy (anonymous):

oh yeah

Parth (parthkohli):

and we didn't do \(29 - 3\) for the range because our list of three numbers is \(5,12,29\)

Parth (parthkohli):

If we had picked \(3\) and \(29\), then the median would have been \(13\) which is not the member of our set... :-P

OpenStudy (anonymous):

yeah i get it

Parth (parthkohli):

;-)

OpenStudy (anonymous):

thnkx